4^{th} Conference

GEOMETRY AND TOPOLOGY OF MANIFOLDS

Krynica 29 IV – 04 V 2002

http://im0.p.lodz.pl/konferencje/krynica2002

Organized by

– Institute of Mathematics of the Technical University of Lodz Address: Al. Politechniki 11, 90-924 Łódź Jan Kubarski (chairman), e-mail: kubarski@ck-sg.p.lodz.pl http://im0.p.lodz.pl/˜ kubarski

– Faculty of Applied Mathematics of the University Mining and Metallurgy, Cracow Adress: al. Mickiewicza 30, 30-059 Kraków, Tomasz Rybicki, e-mail: tomasz@uci.agh.edu.pl http://wms.mat.agh.edu.pl/˜ zrr/

– Institute of Mathematics of the Jagiellonian University, Cracow Address: ul. Wł. Reymonta 4, 30-059 Kraków Robert Wolak, e-mail: wolak@im.uj.edu.pl http://www.im.uj.edu.pl/homepages/Wolak Robert/

Scientific Committee: - - Alekseevsky, Dmitri (Hull, GB), - Belko, Ivan (Minsk, Belarus), - Brown, Ronald (Bangor, UK), - Brzychczy, Stanisław (Cracow, Poland), - Cariñena, Jose (Zaragoza, Spain), - Fernandes, Rui Loja (Lisbon, Portugal), - Grabowski, Janusz (Warsaw, Poland), - Huebschmann, Johannes (Lille, France), - Korbaš, Július (Bratislava, Slovakia), - Kowalski, Oldřich (Prague, Czech Republic), - Kubarski, Jan (Łódź, Poland), - Mishchenko, Alexander (Moscow, Russia), - Nagy, Peter (Debrecen, Hungary), - Nikitin, Anatoly (Kiev, Ukraine), - Pradines, Jean (Toulouse, France), - Rybicki, Tomasz (Cracow, Poland), - Szenthe, János (Budapest, Hungary), - Tulczyjew, Włodzimierz (Camerino, Italy), - Wolak, Robert (Cracow, Poland).

Further information available on our WWW pages:

http://im0.p.lodz.pl/konferencje/krynica2002

SECOND ANNOUNCEMENT

We have the pleasure to inform you that the planned 4th Conference GE-OMETRY AND TOPOLOGY OF MANIFOLDS has met with some interest. The list of potential participants you can find on our WWW page.

Participants of the conference will be lodged in Hotel ”Baśka” of the University of Mining and Metallurgy. We are going to arrange our lectures in two blocks. The first will include longer plenary lectures, which will take place in the morning. The second one, with shorter lectures will be held in the afternoon.

The full day excursion in mountains is planed.

As during the previous meetings, we would like to distribute abstracts of lectures. We kindly ask to send your abstracts (LATEX in the article style or in postscript is preferred) by 10th April 2002. We count on longer ones, even of several pages, or summaries. They will be put on our WWW page. We plan to publish the Proceedings in the Journal ”Universitatis Iagellonicae Acta Mathematica”.

If you have not sent yet the full of Registration Form including the title of your lecture please do it in the first convenience. You can use the Registration Form on our WWW pages.

Full cost of the conference (accommodation and full board) is 200 US $ ( for the accompanying persons 180 US $). The participants are requested to pay at the beginning of the conference.

28 IV (the Sunday) is planed as the day of arrival. The bus of the Jagiellonian University will leave Kraków about the 3 p.m. The departure from Krynica by bus Krynica-Kraków 4 V in the afternoon. The lectures will be held from 29 IV to 4 V 2002. There is a possibility to reserve a hotel room in Cracow for 27/28 IV or 4/5 V. Please inform us about it as soon as possible.

The third announcement including details of the journey to Krynica will be sent in April.

Best regards, organisers:

Jan Kubarski, Tomasz Rybicki, Robert Wolak REGISTRATION FORM:

4^{th} Conference

GEOMETRY AND TOPOLOGY OF MANIFOLDS Krynica 29 IV – 04 V 2002

First and family name: e-mail: http:// Institution: Address of the Institution: Address for correspondence (if different):

Date of arrival: Date of departure:

LECTURE (YES NO) Topic:

COMMUNICATION (YES NO) Topic:

The additional informations: WEGETARIAN (YES NO)

SMOKER (YES NO)

I need a formal invitation (YES NO)

– Institute of Mathematics Technical University of Łódź Al.Politechniki 11, 90-924 Łódź,

– Faculty of Applied Mathematics of the University Mining and Metallurgy, al. Mickiewicza 30, 30-059 Kraków,

– Institute of Mathematics of the Jagiellonian University ul. Wł. Reymonta 4, 30-059 Kraków e-mail: kubarski@ck-sg.p.lodz.pl